# Information-Driven Design of Imaging Systems

### v4

@@ 4 Information estimation on imaging applications
We compared three designs: the traditional Bayer pattern, a random arrangement of red, green, blue and white filters as a baseline, and a learned filter created using the method proposed in [53]. Using the Gehler-Shi dataset of natural images [54, 55], we simulated photon shot noise-limited measurements for each filter design and reconstructed full-color images using a neural network decoder for demosaicing [53]. Information content strongly predicted reconstruction quality across all designs; higher mutual information correlated with better image reconstructions. This held true for multiple image quality metrics: mean squared error (MSE), peak signal-to-noise ratio (PSNR), and structural similarity (SSIM) (Fig. S20). This suggests that different filter designs can be compared using information content alone, avoiding the computational burden of the reconstruction algorithm and the need for ground truth data.
Figure 2: Four imaging applications demonstrating that our mutual information metric predicts downstream performance of system designs. a) Color filter pattern designs in digital cameras: three example scenes from the object distribution, three different possible encoders and the noisy measurements they produce are shown. We estimated mutual information from the measurements only. We also performed decoding (using a neural network) and calculated mean squared error (MSE) from the ground truth, finding that it correlated with our information estimates. b) Potential radio telescope array configurations for black hole imaging. Here, we evaluated three different telescope placements for the Event Horizon Telescope and show that our information estimator predicted the accuracy of recovered black hole structures after decoding by an inverse problem solver. c) Optical elements in lensless cameras, evaluated at different signal levels by the error in deconvolved images. d) LED array illumination patterns for cellular imaging, tested across multiple illumination intensities and evaluated by accuracy of phenotype predictions. In all these examples, information estimation consistently predicted system performance (Fig. S20), validating its use for evaluating imaging systems without requiring decoders or ground truth data.
Black hole imaging
Achieving the necessary angular resolution to image a black hole would require a telescope the size of Earth. While building such a telescope is impossible, astronomers can achieve similar resolution by combining data from radio telescopes across the globe, as demonstrated by the Event Horizon Telescope’s landmark image of the M87 black hole [56]. The next-generation Event Horizon Telescope aims to go beyond static images to create high-fidelity videos of black holes by adding new radio telescopes at carefully chosen sites around the world [57]. However, determining which telescope sites will produce the best images is challenging. The value of each site depends on its relationship to all others, and while measurements from different configurations can be simulated, performing image reconstructions to evaluate every option is computationally prohibitive. Current approaches therefore rely on simplified metrics that do not account for all features of interest [57].
@@
We tested information estimation on a simplified version of the site selection problem (Fig. 2b), selecting combinations of four telescope locations from the original Event Horizon Telescope array. For each configuration, we simulated measurements with additive Gaussian noise and compared our information estimates with quality metrics of the reconstructed image after solving the inverse problem. Measurements with more information consistently produced more accurate reconstructed images of the black hole structure (Fig. S20). Information estimation can thus provide an alternative way to evaluate telescope placement that adapts to the data, without requiring complex reconstruction algorithms.
Lensless imaging
@@
Lensless imagers replace traditional lenses with light-modulating masks, offering compelling advantages: simple hardware, wide field-of-view, and the ability to encode multiple dimensions like depth and time in a single measurement [58]. However, these devices produce encoded measurements that look nothing like the original scene and can vary dramatically between designs, requiring computational reconstruction to recover recognizable images and evaluate performance.
@@
We compared three designs: a traditional lens, a random microlens array [59], and a Gaussian diffuser [60]. Using simulated photon shot noise-limited measurements of natural images at various light levels, we found that information content strongly predicted reconstruction quality across all designs and noise levels (Fig. 2b) (Fig. S20), demonstrating that information estimation can be used to assess quality without relying on reconstructed images.
Coded illumination microscopy
LED array microscopy swaps a microscope’s standard lamp with a programmable array of light sources [61, 62]. Images with different types of contrast can be captured by illuminating the sample with LED array patterns that vary the illumination angles, intensity, and spatial coherence. Evaluating these patterns through classical optical theory is challenging because the optimal patterns will depend greatly on the task at hand. Typically, various LED array patterns are compared empirically [63, 64] by evaluating results after the decoding process for a specific task. We choose here the task of predicting protein expression. Using a dataset of white blood cells [65], we compared three illumination patterns: Brightfield, Differential Phase Contrast [66, 67], and Single LED illumination. Measurement contrast correlates to how cellular structures absorb and scatter light, providing label-free measurements of phenotypic features that correlate with protein expression. To study how noise affects system performance, we added simulated photon shot noise to experimental measurements and evaluated both information content and the accuracy of neural networks in predicting the protein level of individual cells.
Mutual information showed a positive correlation with prediction accuracy across all conditions, though single-LED illumination showed higher information content relative to its prediction performance (likely in part due to its sensitivity to optical imperfections inflating information estimates (Sec. S4.1)). The ability to predict system performance from information content is particularly significant here because measuring information requires only raw images, while traditional evaluation requires both time-consuming protein labeling experiments to generate ground truth datasets, and decoder training.
@@ 4.1 Task-specific information
The microscopy example highlights a key limitation: not all information contributes equally to specific tasks (Sec. S1.5.2). While our method estimates total object information in measurements, task-specific applications like classification may only require certain features. Systems that preserve specific task-relevant information can outperform those capturing more total information indiscriminately [23, 68]. This explains why different illumination patterns showed varying correlations between total information and protein prediction accuracy, despite information consistently increasing with signal-to-noise ratio.
We observed this as well in other contexts. For example, in a 10-class object classification task using lensless imaging (Fig. S22), different optical designs showed varying classification performance despite similar total information content. While classification accuracy improved with signal-to-noise ratio for each design, the total information measured (100$\sim$ 100$\times$ 100 × the amount needed for perfect classification (Sec. S4.2)) did not predict performance across designs. Like other global metrics such as image reconstruction error, mutual information may not always directly correlate with information relevant to specialized tasks.
Information-Driven Encoder Analysis Learning (IDEAL)
Beyond evaluating existing systems, information content can guide the design of new ones. We propose a design framework termed Information-Driven Encoder Analysis Learning (IDEAL), which optimizes imaging systems to maximize the information they capture. IDEAL requires two key components: a differentiable model of the encoder and either a dataset or probabilistic model of the objects being imaged. Because the entropy terms used in our information estimator are also differentiable [69], we can use gradient-based optimization to search for encoder designs that maximize information content (Fig. 3a).
We tested IDEAL for designing an improved color filter array for photography (Fig. 2a), optimizing the pattern of red, green, blue and white filters to maximize information content. Starting from a random configuration, IDEAL progressively improved the design (Fig. 3b). To validate these improvements, we performed two analyses. First, since the optimization objective used the faster but less accurate Gaussian model, we used our most accurate (PixelCNN) estimator to verify that information content increased throughout optimization. Second, we trained image reconstruction algorithms on measurements from different stages of the optimization, finding that the increases in information content corresponded to improvements in reconstruction quality (Fig. 3c).
Figure 3: Information-Driven Encoder Analysis Learning (IDEAL) enables automated design of imaging systems. a) IDEAL combines a differentiable encoder model with our information estimator to optimize system parameters using gradient-based optimization. Here we demonstrate IDEAL on color filter array design for photography using the our Gaussian mutual information estimator as an objective. b) (Gaussian) information estimates increase over the course of optimization as IDEAL refines the color filter pattern. c) Independent validation using the more accurate PixelCNN estimator and reconstruction error confirms that IDEAL’s designs progressively improve both information content and downstream performance.
@@ 5 Discussion
In this work, we developed a practical method for estimating information content in imaging systems using only measurement data and noise characterization. Our approach provides a unified metric that quantifies complex trade-offs between resolution, noise, and other design parameters, as demonstrated across diverse applications from photography to astronomy. Since our estimator requires neither ground truth data nor restrictive assumptions about objects, it can be applied to a wide range of imaging systems. Furthermore, it enables automated, gradient-based optimization of imaging systems designs to maximize information capture.
Our methods address a critical challenge in modern imaging: objectively evaluating systems that rely on computational processing. While algorithms can extract useful information from raw data that is uninterpretable to human observers, evaluating the quality of the measurements is difficult. In the absence of ground truth data, subjective visual assessment of reconstructed images is unreliable, as neural networks can hallucinate convincing but incorrect details [70]. However, ground truth comparisons across multiple experimental systems remain impractical or impossible. By quantifying information content directly from measurements, our approach offers a way to objectively evaluate and compare real-world instruments.
While our approach enables finding optimal data capture methods, it does not solve the separate challenge of decoding measurements into a useful form. High information content indicates potential for excellent performance, but realizing this potential still requires developing appropriate algorithms to extract and interpret the encoded information.
@@
The ability to quantify information may reveal that characteristics of measurements traditionally considered noise actually contain useful signal. Information theory suggests that optimal encoders would produce random-looking measurements to maximize entropy and ensure small object changes remain distinguishable [71, 72]. Such measurements would appear noisy and unintuitive to human observers, who are selectively sensitive to specific visual features [1]. However, it may be possible to extract this signal computationally. For example, laser speckle patterns have traditionally been considered noise and suppressed [32], but more recent work shows they can be effectively interpreted by neural networks [73, 74, 75], suggesting that the useful signal they contain is merely unintelligible to human observers.
Future work could advance the accuracy-compute-data Pareto frontier of our estimators. Since our approach reduces to density modeling, it could leverage transformer architectures [76] that exhibit predictable scaling laws [44, 77]. This enables either deploying larger models for more accurate direct estimates or extrapolating the true entropy using scaling relationships on smaller models. Alternatively, specialized architectures might achieve equivalent accuracy with fewer resources. Validating improvements remains straightforward: any model achieving lower cross-entropy on fixed data necessarily tightens the bound on true information content.
Our Information-Driven Encoder Assessment Learning (IDEAL) pipeline shows promising initial results that suggest key computational advantages over existing “end-to-end” optimization methods [78, 79, 80, 81, 82, 83, 84]. Unlike end-to-end methods that simultaneously optimize both the imaging system (encoder) and reconstruction algorithm (decoder), IDEAL optimizes only the encoder, allowing decoders to be designed or optimized later. This significantly reduces computation and memory requirements and avoids challenges associated with backpropagating through complex neural network decoders [84, 85], such as vanishing gradients.
IDEAL’s optimization of general information capture, rather than task-specific performance, presents both advantages and limitations. Our experiments show that while general information content correlates with system performance, this relationship weakens for specialized applications that depend on particular object features. Task-specific methods may achieve better performance in these cases, though they typically require specialized training data or specific assumptions about the objects being imaged [23, 68, 86, 85]. When such assumptions or data are unavailable, IDEAL provides a robust default strategy. IDEAL-optimized systems could also serve as effective starting points for task-specific refinement, similar to how machine learning models are pre-trained on large unlabeled datasets before specialized fine-tuning. This hybrid approach would avoid the computational challenges of full task-specific optimization while still enabling specialized performance with limited additional data.
In summary, our information-based approach offers a versatile tool for evaluating and designing diverse imaging systems. The method’s generic probabilistic foundation suggests potential applications beyond imaging, extending to electronic, biological, geological, and chemical sensors.
@@ Acknowledgements
This section assumes familiarity with probability and information theory fundamentals. A tutorial introduction to these concepts can be found in previous work [87], as well as the textbooks [69, 72] and Shannon’s original paper [71].
@@
The minimal probabilistic model we present has been widely used in information-theoretic analyses of imaging systems [5, 14, 22, 31] and builds on established principles in statistical optics [32].
@@
  • Forward: Object → Noiseless Image → Noisy Measurement ($p(\mathbf{o},\mathbf{x},\mathbf{y})=p(\mathbf{o})p(\mathbf{x}|\mathbf{o})p(\mathbf{y}|\mathbf{x})$)
  • Reverse: Noisy Measurement → Noiseless Image → Object ($p(\mathbf{o},\mathbf{x},\mathbf{y})=p(\mathbf{y})p(\mathbf{x}|\mathbf{y})p(\mathbf{o}|\mathbf{x})$)
The Data Processing Inequality [72] reveals key limitations in each direction:
  • Forward: $I(\mathbf{O};\mathbf{X})\geq I(\mathbf{O};\mathbf{Y})$ noise can only reduce information about the object.
  • Reverse: $I(\mathbf{Y};\mathbf{X})\geq I(\mathbf{Y};\mathbf{O})$ the measurement’s ability to carry information about the object is limited by the information it carries about the noiseless image.
@@
This minimal model, which separates image encoding from detection noise, has been widely used in information-theoretic analyses of imaging systems [5, 14, 22, 31] and general imaging models [88]. For optical systems, this separation is justified by the “semi-classical” approach common in statistical optics [32]. This method treats light classically until it reaches the detector, where quantum effects introduce noise. Optical systems can experience two types of noise:
@@
The probability distributions over images ($p(\mathbf{y})$) and measurements ($p(\mathbf{x})$) are high-dimensional, and thus not possible to visualize directly. In microscopy, for example (Fig. S2a), a microscope transforms a distribution of possible cells into noiseless images, which detection then converts into noisy measurements. To provide insight into these complex distributions, we adapt a representation from [89] (Fig. S2b). This representation, which we call energy coordinates, offers a complementary perspective to the traditional spatial coordinate representation, enabling a more comprehensive understanding of image distributions and information flow in imaging systems.
@@
With any model, the space of noiseless images is ultimately finite in some sense. Energy cannot be infinitely concentrated in a single point111…without collapsing space into a black hole [90], and the physics of wave-propagation effectively constrain electromagnetic waves to a finite number of degrees of freedom [90]. This means that even in the continuous/continuous case, any noiseless image can be represented to an arbitrary level of precision by a finite number of samples.
@@
We demonstrate with a simplified scenario, where the task is to distinguish between one point source or two closely spaced dimmer point sources, in the presence of measurement noise. We can derive the information content analytically for this idealized case (see Section S1.4.1) in which we model the object as either one point or two half-energy points, with equal probability [91, 92]. We assume that the imaging system has a diffraction-limited point spread function, with measurements corrupted by additive Gaussian noise (Fig. S3a). This simplified setup allow us to examine how information content accounts for both resolution and signal-to-noise ratio.
@@
The family of encoders $\mathcal{E}$ studied were 1D bandlimited, nonnegative, linear-shift invariant, infinitely periodic point spread functions. Figure S4a shows the outputs of a representative encoder in this family (i.e. a specific point spread function) acting on a distribution of delta function objects. This system can be thought of as a simplified version of an imaging system that uses spatially incoherent illumination, such as in photography or microscopy [93].
@@
The set of possible output signals $\mathcal{\boldsymbol{X}}$ for this family of encoders is identical to the set of possible point spread functions: bandlimited, nonnegative, and infinitely periodic signals (Fig. S6b). These output signals (which are analogous to the noiseless images in discussed in Formulation of information estimator) can be viewed either in spatial coordinates, in which their energy density is plotted as a function of space, or in energy coordinates (Fig. S2), with values found by integrating areas of the signal corresponding to “pixels.”
Signals cannot have both finite bandwidth and finite extent in space, and they thus require infinite samples to represent with arbitrary accuracy in the absence of further constraints. This is typically handled in one of two ways: treating signals as both band-limited and spatially finite and sampling at finite density over finite extent; or considering band-limited signals with infinite spatial extent, which asymptotically have finite degrees of freedom and can thus be sampled at finite rates with arbitrary accuracy [90].
@@
As a result of this simplification, there is a bijective mapping between the set of possible signals and non-negative vectors in $D$-dimensional space–that is, each signal in the set corresponds to a point in $D$-dimensional space (similar to the argument made in [89]). This allows us to computationally analyze this finite-dimensional space with insights that can be applied to the more complicated space of continuous signals.

### v5

@@ 4 Validation across imaging applications
Our information estimates predicted reconstruction quality across three color filter designs, enabling the evaluation of designs without reconstruction algorithms or ground truth data (Fig. 2a). We tested the traditional Bayer pattern, a random arrangement of red, green, blue, and white filters, and a learned arrangement of these filters [chakrabartiLearningSensorMultiplexing]. Using natural images [gehlerBayesianColorConstancy2008a, shiReprocessedVersionGehler] with simulated photon shot noise, we estimated information content for each design and reconstructed full-color images using neural network demosaicing [chakrabartiLearningSensorMultiplexing]. Higher mutual information consistently predicted better reconstruction across all quality metrics—mean squared error (MSE), peak signal-to-noise ratio (PSNR), and structural similarity index measure (SSIM) (Fig. S20).
Figure 2: Information estimates predict decoder performance across four imaging applications. Each row shows representative objects, three encoder designs, example measurements, and the relationship between information estimates and a representative decoder performance metric (see Fig. S20 for all metrics). a) Color photography: Bayer, learned, and random filter arrays with neural network demosaicing [chakrabartiLearningSensorMultiplexing]. b) Radio astronomy: Three telescope array configurations with inverse problem reconstruction. c) Lensless imaging: Lens, microlens array, and diffuser with Wiener deconvolution. d) LED array microscopy: Brightfield, differential phase contrast, and single-LED illumination with neural network protein expression prediction.
Black hole imaging.
Radio telescope arrays can achieve the angular resolution of an Earth-sized telescope by combining data from global sites, enabling images like the Event Horizon Telescope’s M87 black hole image [theeventhorizontelescopecollaborationFirstM87Event2019]. Selecting optimal telescope locations for next-generation arrays remains computationally challenging: each site’s value depends on all others, and reconstructing images for every possible configuration is computationally intractable [doelemanReferenceArrayDesign2023]. Thus, current approaches use simplified metrics that can miss important imaging features [doelemanReferenceArrayDesign2023].
@@
Our information estimates predicted reconstruction quality across telescope array configurations, enabling telescope site selection without computationally expensive image reconstruction (Fig. 2b). Using four-telescope subsets from the original Event Horizon array, we simulated measurements with additive Gaussian noise and compared information estimates to the error in black hole images generated by an iterative inverse problem optimization. Higher information estimates consistently produced more accurate reconstructions (Fig. S20).
Lensless imaging.
@@
Lensless imagers replace traditional lenses with light-modulating masks, offering simple hardware, wide field-of-view, and the ability to encode depth and time information in single measurements [boominathanRecentAdvancesLensless2022]. Their encoded measurements bear no visual resemblance to scenes and vary dramatically between designs, requiring computational reconstruction for both image recovery and performance evaluation.
@@
Our information estimates predicted reconstruction quality across three optical designs, enabling evaluation of unconventional optics without image reconstruction (Fig. 2c). We tested a traditional lens, random microlens array [kabuliHighQualityLenslessImaging2023], and Gaussian diffuser [antipaDiffuserCamLenslessSingleexposure2018] using natural images with simulated photon shot noise at various light levels. Higher information estimates consistently correlated with better reconstruction accuracy across all designs and noise conditions (Fig. S20).
Coded illumination microscopy.
LED array microscopy replaces standard illumination with programmable light sources to flexibly generate different contrast modes at low cost by varying angle, intensity, and coherence [zhengMicroscopyRefocusingDarkfield2011, phillipsQuasiDomeSelfCalibratedHighNA2017]. Optimal illumination patterns depend heavily on imaging tasks, making theoretical evaluation difficult and necessitating empirical comparison through task-specific decoders [kellmanPhysicsBasedLearnedDesign2019, horstmeyerConvolutionalNeuralNetworks2017].
Our information estimates correlated with protein prediction accuracy across three illumination patterns, enabling evaluation of microscopy designs without time-consuming and expensive protein labeling experiments (Fig. 2d). We tested brightfield, differential phase contrast [mehtaQuantitativePhasegradientImaging2009, tianQuantitativeDifferentialPhase2015], and single-LED illumination measurements of white blood cells [pinkardBerkeleySingleCell2024] with added simulated photon shot noise to equalize photon counts. Higher information estimates generally predicted better neural network classification performance, though single-LED showed inflated estimates due to optical imperfections (Sec. S4.1).
@@ 4.1 Task-specific information
While information estimates strongly predict reconstruction performance, they may correlate less strongly with specialized tasks that rely only on specific features of measurements. Systems that preserve specific task-relevant information can outperform those capturing more total information indiscriminately [neifeldTaskspecificInformationImaging2007, ashokCompressiveImagingSystem2008].
Our experiments confirmed that specialized tasks show weaker correlations between information estimates and decoder performance. We tested this by switching from reconstruction to a classification task for the lensless imaging system (Fig. S22). This classification task uses only $\sim$1% of the information in measurements (Sec. S4.2). We found that designs with similar total information achieved different classification accuracies, likely reflecting different amounts of task-specific information captured (Fig. S20). A similar pattern may explain the microscopy results (Fig. 2d), where single-LED illumination yielded less accurate protein predictions than differential phase contrast despite higher information estimates.
These findings motivate extensions that optimize task-specific rather than total information. By decomposing information into task-specific and task-irrelevant components ($I(\mathbf{Y};\mathbf{O})=I(\mathbf{Y};\mathbf{T})+I(\mathbf{Y};\mathbf{O}|\mathbf{T})$), future work could target specialized tasks while maintaining decoder-independent evaluation (Sec. S2.8).
5 Encoder design via information maximization with IDEAL
Having validated that information estimates can assess encoder quality directly from measurements, we tested their ability to guide computationally efficient design optimization. To realize this capability, we developed Information-Driven Encoder Analysis Learning (IDEAL), which iteratively improves encoder designs through gradient ascent on information estimates (Fig. 3a). IDEAL requires only a differentiable encoder model and a dataset or model representing the objects to be imaged, avoiding the memory and compute requirements and backpropagation challenges of complex neural network decoders.
IDEAL successfully optimized color filter arrays, creating encoders that capture more information and enable better reconstruction. Testing on photography filters with red, green, blue and white pixels, IDEAL progressively improved filter designs compared to a random initial filter (Fig. 3b). To validate the optimization, we trained decoders on measurements from designs at different optimization stages, and found that reconstruction accuracy increased from beginning to end (Fig. 3c). Furthermore, our PixelCNN estimator independently confirmed that information content increased throughout optimization, validating that the faster Gaussian model used as our objective successfully guided the design process. IDEAL achieved comparable information content and decoder error to end-to-end optimization [chakrabartiLearningSensorMultiplexing], which jointly trains both encoders and decoders. This encoder-only approach offers advantages in simplicity, memory, and runtime, which are studied in greater depth in subsequent work [markleyComputationallyEfficientInformationDriven2025].
Figure 3: Information-Driven Encoder Analysis Learning (IDEAL) designs encoders through gradient ascent on information estimates. a) IDEAL framework applied to color filter design: gradient feedback from information estimates updates parameters of a differentiable filter model to maximize information capture. b) Estimated information increases monotonically during optimization. c) The final IDEAL-optimized encoder matches the performance of jointly optimizing encoders and decoders end-to-end [chakrabartiLearningSensorMultiplexing] in terms of both downstream reconstruction error and measurement mutual information, while avoiding decoder complexity during training. We further study IDEAL in comparison to end-to-end in subsequent work [markleyComputationallyEfficientInformationDriven2025].
@@ 6 Discussion
In this work, we developed a broadly applicable and computationally efficient method that comprehensively captures imaging system performance by quantifying the information content of measurements. The method applies broadly because it uses only measurement data and noise characterization, without requiring any knowledge or assumptions about objects or system-specific mathematical models. It captures comprehensive performance because mutual information naturally synthesizes resolution, noise, sampling, and all other factors that affect the ability to distinguish different objects into a single metric. It achieves computational efficiency by evaluating encoders directly, bypassing the traditional approach of using decoder algorithms to perform secondary tasks and then measuring success on those tasks.
We validated that our approach realizes these advantages while matching traditional decoder-based evaluation and design strategies across color photography, radio astronomy, lensless imaging, and microscopy. Information estimates consistently predicted which designs would enable decoders to reconstruct unknown objects most accurately. Additionally, we demonstrated that information estimates can guide automated design of imaging systems using IDEAL (Information-Driven Encoder Analysis Learning), achieving comparable performance to the prevailing approach of end-to-end optimization through an encoder and decoder simultaneously.
Information-based evaluation offers computational advantages that suggest promising capabilities for designing previously intractable imaging systems in simulation. By avoiding decoders, the approach eliminates the memory requirements of large neural networks and the training complexity of joint optimization, such as vanishing gradients that arise during backpropagation through complex architectures [debFourierNetsEnableDesign2022, yangImageQualityNot2023a]. This reduced computational burden enables more effective exploration of design spaces and may allow for tackling new problems where end-to-end design becomes computationally intractable. We explore IDEAL’s capabilities and limitations more extensively across diverse imaging systems in subsequent work [markleyComputationallyEfficientInformationDriven2025].
Beyond simulated design, our method creates entirely new possibilities for rigorously evaluating imaging systems operating in real-world conditions. Current evaluation approaches each have significant limitations: subjective visual assessment of raw measurements or algorithm outputs suffers from observer bias and inconsistency; heuristic quality metrics applied to algorithm outputs are unreliable for powerful neural network decoders that can hallucinate convincing but incorrect details [jiSurveyHallucinationNatural2023]; assessments like signal-to-noise ratio capture only isolated system aspects rather than overall capability; and standardized test objects like resolution targets may not represent the diversity of real-world objects. Our information-based approach addresses these limitations by providing an objective, unified metric that quantifies measurement quality without requiring knowledge or assumptions about the objects being imaged. Our approach thus enables rigorous and comprehensive evaluation of imaging systems capturing unknown objects in their intended environments.
@@
More broadly, our method may aid in discovering signal in measurements that appear unintelligible to human observers. Information theory suggests that optimal encoders produce measurements that appear random or noisy to maximize entropy and ensure small object changes remain distinguishable [shannonMathematicalTheoryCommunication1948, coverElementsInformationTheory2005]. Such measurements would naturally appear unintuitive to human observers who are selectively sensitive to specific visual features [fassHumanSensitivityMutual]. Laser speckle patterns hint at this possibility: once routinely suppressed as unwanted noise[goodmanStatisticalOptics2015], they can now be effectively interpreted by neural networks[mirzaalianEffectivenessLaserSpeckle2019, liDeepSpeckleCorrelation2018, kurumDeepLearningEnabled2019].
Extensions and future directions
Beyond our framework’s primary applications to the evaluation and design of imaging systems, several promising directions could improve its capabilities, extend its theoretical foundations, and enable its application to new classes of design problems.
Performance improvements. Better probabilistic models could improve the estimator’s accuracy. Transformer architectures [vaswaniAttentionAllYou2017] exhibit predictable scaling laws [henighanScalingLawsAutoregressive2020, openaiGPT4TechnicalReport2024], enabling either deployment of larger models for more accurate estimates or extrapolation of true information content using scaling relationships on smaller models. Alternatively, specialized architectures might achieve equivalent accuracy with fewer computational resources.
Stochastic encoders. Extending the framework to handle non-deterministic imaging systems would broaden its applicability. Two scenarios require this extension: systems with unpredictable variations like illumination fluctuations or mechanical drift, and quantum imaging regimes where image formation is inherently random [higginsEntanglementfreeHeisenberglimitedPhase2007, giovannettiQuantumEnhancedMeasurementsBeating2004, israelSupersensitivePolarizationMicroscopy2014]. We outline key challenges for this extension in the supplement (Sec. S2.7), though practical validation remains future work.
Task-specific imaging. Extending the framework to specialized tasks would address applications where maximizing total information may be suboptimal. While our approach captures performance well on reconstruction tasks that require recovering objects in complete detail, information estimates less accurately characterize performance when decoders perform tasks such as classification that prioritize specific object features. Modifying the objective function our estimator  (Sec. S2.8) might address this limitation. Alternatively, optimization for general performance with IDEAL could provide a starting point for task-specific refinement using end-to-end design methods [yangImageQualityNot2023a] that reduces labeled data requirements.
Design in non-differentiable systems. Information estimation could enable optimization of non-differentiable imaging systems where neither end-to-end design nor IDEAL would apply. This includes combinatorial design problems in simulation and real-time parameter adjustments in deployed imaging systems. Both applications require efficient objective functions that can be evaluated repeatedly during gradient-free optimization. Subsequent work has applied information estimation to non-differentiable imaging system design [kabuli2025blackboxinfo].
Other sensing modalities. Information-theoretic design principles could extend beyond imaging to optimize performance in electronic, biological, geological, and chemical sensors. Our method could potentially apply to any sensing system that can be modeled as a deterministic encoding with a known noise model, suggesting possibilities for systematic sensor design across diverse domains.
@@ Acknowledgements
This section assumes familiarity with probability and information theory fundamentals. A tutorial introduction to these concepts can be found in previous work [pinkardVisualIntroductionInformation2022], as well as the textbooks [mackayInformationTheoryInference2003, coverElementsInformationTheory2005] and Shannon’s original paper [shannonMathematicalTheoryCommunication1948].
@@
The minimal probabilistic model we present has been widely used in information-theoretic analyses of imaging systems [fellgettAssessmentOpticalImages1955, huckInformationTheoryVisual1997, ashokInformationbasedAnalysisSimple2003, changInformativeSensing2009] and builds on established principles in statistical optics [goodmanStatisticalOptics2015].
@@
  • Forward: Object → Noiseless Image → Noisy Measurement
    $p(\mathbf{o},\mathbf{x},\mathbf{y})=p(\mathbf{o})p(\mathbf{x}|\mathbf{o})p(\mathbf{y}|\mathbf{x})$
  • Reverse: Noisy Measurement → Noiseless Image → Object
    $p(\mathbf{o},\mathbf{x},\mathbf{y})=p(\mathbf{y})p(\mathbf{x}|\mathbf{y})p(\mathbf{o}|\mathbf{x})$
The Data Processing Inequality [coverElementsInformationTheory2005] reveals key limitations in each direction:
  • Forward: $I(\mathbf{O};\mathbf{X})\geq I(\mathbf{O};\mathbf{Y})$ - noise can only reduce information about the object.
  • Reverse: $I(\mathbf{Y};\mathbf{X})\geq I(\mathbf{Y};\mathbf{O})$ - the measurement’s ability to carry information about the object is limited by the information it carries about the noiseless image.
@@
Deterministic Information Transfer Theorem. For a Markov chain $\mathbf{O}\to\mathbf{X}\to\mathbf{Y}$, if $\mathbf{X}=f(\mathbf{O})$ for some deterministic function $f$, then $I(\mathbf{O};\mathbf{Y})=I(\mathbf{X};\mathbf{Y})$.
Proof.
Since $I(\mathbf{O};\mathbf{Y})=H(\mathbf{Y})-H(\mathbf{Y}|\mathbf{O})$ and $I(\mathbf{X};\mathbf{Y})=H(\mathbf{Y})-H(\mathbf{Y}|\mathbf{X})$, we need to show $H(\mathbf{Y}|\mathbf{O})=H(\mathbf{Y}|\mathbf{X})$.
From the deterministic condition $\mathbf{X}=f(\mathbf{O})$: knowing $\mathbf{O}$ completely determines $\mathbf{X}$. Therefore, conditioning on $\mathbf{O}$ is equivalent to conditioning on both $\mathbf{O}$ and $\mathbf{X}$:
$H(\mathbf{Y}|\mathbf{O})=H(\mathbf{Y}|\mathbf{O},\mathbf{X})$ (5)
From the Markov property $\mathbf{O}\to\mathbf{X}\to\mathbf{Y}$:
$H(\mathbf{Y}|\mathbf{O},\mathbf{X})=H(\mathbf{Y}|\mathbf{X})$ (6)
Therefore, $H(\mathbf{Y}|\mathbf{O})=H(\mathbf{Y}|\mathbf{X})$, which implies $I(\mathbf{O};\mathbf{Y})=I(\mathbf{X};\mathbf{Y})$. ∎
@@
This minimal model, which separates image encoding from detection noise, has been widely used in information-theoretic analyses of imaging systems [fellgettAssessmentOpticalImages1955, huckInformationTheoryVisual1997, ashokInformationbasedAnalysisSimple2003, changInformativeSensing2009] and general imaging models [ronchiResolvingPowerCalculated1961]. For optical systems, this separation is justified by the “semi-classical” approach common in statistical optics [goodmanStatisticalOptics2015]. This method treats light classically until it reaches the detector, where quantum effects introduce noise. Optical systems can experience two types of noise:
@@
The probability distributions over images ($p(\mathbf{y})$) and measurements ($p(\mathbf{x})$) are high-dimensional, and thus not possible to visualize directly. In microscopy, for example (Fig. S2a), a microscope transforms a distribution of possible cells into noiseless images, which detection then converts into noisy measurements. To provide insight into these complex distributions, we adapt a representation from [shannonCommunicationPresenceNoise1949] (Fig. S2b). This representation, which we call energy coordinates, offers a complementary perspective to the traditional spatial coordinate representation, enabling a more comprehensive understanding of image distributions and information flow in imaging systems.
@@
With any model, the space of noiseless images is ultimately finite in some sense. Energy cannot be infinitely concentrated in a single point111…without collapsing space into a black hole [franceschettiWaveTheoryInformation2017], and the physics of wave-propagation effectively constrain electromagnetic waves to a finite number of degrees of freedom [franceschettiWaveTheoryInformation2017]. This means that even in the continuous/continuous case, any noiseless image can be represented to an arbitrary level of precision by a finite number of samples.
@@
We demonstrate with a simplified scenario, where the task is to distinguish between one point source or two closely spaced dimmer point sources, in the presence of measurement noise. We can derive the information content analytically for this idealized case (see Section S1.4.1) in which we model the object as either one point or two half-energy points, with equal probability [harrisResolvingPowerDecision1964, heintzmannTwoPointResolution2001]. We assume that the imaging system has a diffraction-limited point spread function, with measurements corrupted by additive Gaussian noise (Fig. S3a). This simplified setup allow us to examine how information content accounts for both resolution and signal-to-noise ratio.
@@
The family of encoders $\mathcal{E}$ studied were 1D bandlimited, nonnegative, linear-shift invariant, infinitely periodic point spread functions. Figure S4a shows the outputs of a representative encoder in this family (i.e. a specific point spread function) acting on a distribution of delta function objects. This system can be thought of as a simplified version of an imaging system that uses spatially incoherent illumination, such as in photography or microscopy [goodmanIntroductionFourierOptics2005].
@@
The set of possible output signals $\mathcal{\boldsymbol{X}}$ for this family of encoders is identical to the set of possible point spread functions: bandlimited, nonnegative, and infinitely periodic signals (Fig. S6b). These output signals (which are analogous to the noiseless images in discussed in Mathematical framework) can be viewed either in spatial coordinates, in which their energy density is plotted as a function of space, or in energy coordinates (Fig. S2), with values found by integrating areas of the signal corresponding to “pixels.”
Signals cannot have both finite bandwidth and finite extent in space, and they thus require infinite samples to represent with arbitrary accuracy in the absence of further constraints. This is typically handled in one of two ways: treating signals as both band-limited and spatially finite and sampling at finite density over finite extent; or considering band-limited signals with infinite spatial extent, which asymptotically have finite degrees of freedom and can thus be sampled at finite rates with arbitrary accuracy [franceschettiWaveTheoryInformation2017].
@@
As a result of this simplification, there is a bijective mapping between the set of possible signals and non-negative vectors in $D$-dimensional space–that is, each signal in the set corresponds to a point in $D$-dimensional space (similar to the argument made in [shannonCommunicationPresenceNoise1949]). This allows us to computationally analyze this finite-dimensional space with insights that can be applied to the more complicated space of continuous signals.